OpenStudy (anonymous):

An aircraft flew 3 hours with the wind. The return trip took 4 hours against the wind. If the speed of the plane in still air is 234 miles per hour more than the speed of the wind, find the wind speed and the speed of the plane in still air. a. the wind speed was b. the speed of the plane in still air was

OpenStudy (anonymous):

(S_airplane+S_wind)*3=(S_airplane-S_wind)*4

OpenStudy (anonymous):

Can you give me the answer?? I cant figure it out. ill give you a medal

OpenStudy (anonymous):

airplane in still air =234+S ((234+S)+S)*3=((234+S)-S)*4 3*234+6S=234*4 6S=234 S=234/6=39 mil/hr

OpenStudy (anonymous):

this the speed of the wind 39 mil/hr airplane= 234+39=273 mil/hr

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

Yw

OpenStudy (aaronandyson):

was going to write a full sets of figure explaining this but @AJ01 already did so.. O.O